OpenStudy (anonymous):

Let f(x) = twenty-four over the quantity of 3 x minus two. Find f(-2).

OpenStudy (anonymous):

\[f(x)=\frac{24}{3x-2}\] \[f(-2)=\frac{24}{3\times (-2)-2}\]

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

now what what?

OpenStudy (anonymous):

like what do you do after that.

OpenStudy (anonymous):

compute the number i guess

OpenStudy (anonymous):

your answer is a number it is what you get when you compute \[\frac{24}{3\times (-2)-2}\]